OpenStudy (anonymous):

Describe a pattern in each sequence. What are the next two terms of each sequence? -5, 5, -5, 5, . . . Select one: a. multiply the previous term by 1; 5, 5 b. multiply the previous term by -1; -5, 5 c. add 10 to the previous term; -5, -15 d. subtract 10 from the previous term; -5, 5

OpenStudy (anonymous):

it would be b

OpenStudy (anonymous):

b for sure

OpenStudy (anonymous):

B

OpenStudy (anonymous):

hey can anyone help me with my math prob?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I will help you

OpenStudy (anonymous):

will give a medal! For the remaining three problems, write always, sometimes, or never. If it is always true, give the property or two examples. If it is never true, give the property or two examples. If it is sometimes true, give one example for which it is true and one example for which it is false. One point for a correct answer and one point for examples or correct property. x + x = 2x x • x = 2x x squared = -9 Please help me i will be so thankful!

OpenStudy (anonymous):

its kinda hard tho

OpenStudy (anonymous):

plug in the -9 for the x's and then add and then multiply

OpenStudy (anonymous):

1. Always true ie. 3 + 3 = 2*6 --> 6 = 6 True, 4 + 4 = 2*4 -->8 = 8 True. 2. Sometimes, only two values make it true. ie: 0,2 3. Never true, because when you multiply any two of the same real numbers (negative or positive) together they equal a positive.

OpenStudy (anonymous):